OpenStudy (breann):

2 3/4 divided by 1 1/2= i got 11/4d divided by 3/2= but unsure what to do after that....

OpenStudy (anonymous):

11/4 divided by 3/2 = 11/4 multiplied by 2/3. The rest follows naturally.

OpenStudy (anonymous):

hey moorsy come back to kaya4me I have a question

OpenStudy (hoblos):

[11/4 \div 3/2 = 11/4 \times 2/3 =11/6\]

OpenStudy (breann):

hoblos, how did you get the denominator to be 6?

OpenStudy (hoblos):

 by that you get 11/6